OpenStudy (anonymous):

How would you go about solving for D D= e^-3ln10 I know the answer is 10^-3 but i dont see exactly how, thanks

OpenStudy (anonymous):

\[-3\ln(10)=\ln(10^{-3})\] is a start

OpenStudy (anonymous):

by the property of the log that says \(n\ln(x)=\ln(x^n)\)

OpenStudy (anonymous):

then \[e^{\ln(\spadesuit)}=\spadesuit\] as it is the composition of a function with its inverse

OpenStudy (anonymous):

that is why \[e^{\ln(10^{-3})}=10^{-3}\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw